# Dissipative solutions to Stochastic 3D Euler equations

### v2

@@ 3.5. Estimates on the velocity perturbation
3.5.1. Proof of Proposition˜3.3: Preliminary estimates
We report some estimates that will be helpful in the proof of Proposition˜3.3, as well as in the remainder of the paper.
@@
Lemma˜3.4 and (2.25) imply the following estimates, that will be useful later on.
@@
First of all, we have that
@@
The same computations and Lemma˜3.4 yield the result for $\|D_{t,l}(v_{q}-v_{l})(t)\|_{N}$. ∎
@@
where $(\nabla u_{l})^{2}$ denotes the standard product between matrices $(\nabla u_{l})^{2}]^{i,j}=\sum_{k}\partial_{i}u^{k}\partial_{k}u^{j}$, then by [DLK23, Lemma A.3],Lemma˜3.4, and the bound $L\,i_{q}^{-1/2-\delta}\leq\lambda^{2}_{q}\delta_{q}$ we have:
(3.66) $\|D_{t,l}\nabla u_{l}(t)\|_{N}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}+L\,i_{q}^{-1/2-\delta}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}.$
Then, the estimate (3.64) on $\|D_{t,l}\nabla v_{l}(t)\|_{N}$ follows from the identity $D_{t,l}\nabla v_{l}(t)=D_{t,l}\nabla u_{l}(t)-D_{t,l}\nabla z_{l}(t)$ and computations similar to those above. ∎
3.5.2. Proof of Proposition˜3.3: Estimates on the oscillatory and compressible Fourier coefficients
@@
Let us start with estimates aimed to control the Fourier coefficients appearing $w$ and in the products $w_{o}\otimes w_{o}$ and $|w_{o}|^{2}w_{o}$, cf. Equations (3.24), (3.25), (3.26), (3.33).
@@
For $s\in\{0,1,2\}$ and $N\leq n_{0}+3$ , it holds for all $t\leq\mathfrak{t}$ that
@@
(3.71) $X_{I}(t,x):=\theta_{m}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(x,t)\overline{f}_{I}(x,t)\mathring{b}_{I,k}.$
@@
(3.72) $|D_{t,l}^{s}\theta_{m}(t)|\lesssim\epsilon^{-s}.$
Second, by Proposition˜A.1 and (3.64), together with the assumed bound $\epsilon\|u_{l}(t)\|_{1}<1$ and the algebraic identities
@@
(3.74) $\|D_{t,l}^{s}\overline{f_{I}}(t)\|_{K}\lesssim l^{-K}(\lambda_{q}\delta_{q}^{1/2})^{s},\quad\text{for }s\in\{0,1,2\}.$
Third, by Proposition˜A.1, (3.20) and [BD15+, Proposition C.1], we have
@@
For $s=1$ instead, we have by (3.73), Proposition˜A.1 and Eq.˜3.36:
@@
where $M_{m,n}$ is given by Eq.˜3.44. Then, (3.73), (3.36), (3.84) imply that
(3.86) $\|D_{t,l}^{s}\left((\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\|_{N}\lesssim\epsilon^{-s}\mu^{-N}.$
Let us stress that we have exploited the bound (3.84) since $M_{m,n}(t,x)$ is given by a finite summation of $X_{J}$ for $J\in\mathcal{I}_{\phi}$. Therefore,
@@
3.5.3. Proof of Proposition˜3.3: Conclusion
@@
(3.89) $D_{t,l}^{s}w_{o}=\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}D_{t,l}^{s}b_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
@@
(3.90) $\|e^{i\lambda_{q+1}\xi_{m}(t)\cdot k}\|_{N}\lesssim|k|^{N}\lambda^{N}_{q+1},$
and thus the desired inequality follows from Lemma˜3.8 and the analog of [DLK23, Equation (3.24)].
@@ 4. Construction of the new Reynolds stress and pressure
(4.8) $\operatorname{Tr}\big(R_{q+1}\big)=\operatorname{Tr}\big(R_{q}-R_{l}\big)+|w_{c}|^{2}+2w_{c}\cdot w_{o}+2(v_{q}-v_{l})\cdot w+2f(t).$
@@ 4.1. Estimates on the Reynolds stress
We denote $D_{t,q+1}:=\partial_{t}+u_{q+1}\cdot\nabla$. Our main proposition on the Reynolds stress is the following: morally speaking, taking either a space derivative or and advective derivative of the new Reynolds stress $R_{q+1}$ “costs” a factor $\lambda_{q+1}$.
@@
(4.10) $\left\|D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right)w(t)\right\|_{N-1}$ $\leq\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2},$
@@
(4.12) $\left\|D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\,\right)\right\|_{N-1}$ $\leq\lambda_{q+1}^{N-2\gamma}\delta_{q+2}\delta_{q+1}^{1/2}.$
Before proving the previous proposition, let us introduce the following parameters:
@@
They will play a crucial role when perfoming estimates involving quantities in which we need to apply the antidivergence operator $\mathcal{R}$, more specifically in the bounds (A.19), (A.46), (A.60) from the Appendix. The values of $n_{0},h_{0}$ ensure that
@@
4.1.1. Proof of Proposition˜4.1: bounds on $R_{T},R_{A},R_{S}$.
As a first step towards proving Proposition˜4.1, we separately give bounds on the stress tensors $R_{T},R_{A},R_{S}$ appearing in the decomposition of the new Reynolds stress $R_{q+1}$.
@@
(4.20) $\mathbf{P}_{\leq l^{-1}}R_{q}-R_{l}(t,x)=\int_{-l_{\rm temp}}^{0}ds\,\eta_{l_{\rm temp}}(s)\int_{s}^{0}(D_{t,l}\mathbf{P}_{\leq l^{-1}}R_{q})(t+r,\mu(t;t+r,x))dr.$
@@
(4.23) $\|\mathbf{P}_{\leq l^{-1}}R_{q}(t)-R_{l}(t)\|_{N}\lesssim l^{-N}\lambda_{q}^{1-\gamma}\delta_{q+1}\delta_{q}^{1/2}l_{\rm temp}.$
@@
(4.25) $\|D_{t,q+1}(R_{q}-R_{l})(t)\|_{N-2}$ $\leq\|D_{t,q+1}R_{q}(t)\|_{N-2}+\|D_{t,q+1}R_{l}(t)\|_{N-2}$
$\leq\|D_{t,q}R_{q}(t)\|_{N-2}+\|(u_{q+1}-u_{q})\cdot\nabla R_{q}(t)\|_{N-2}$
$+\|D_{t,l}R_{l}(t)\|_{N-2}+\|(u_{q+1}-u_{l})\cdot\nabla R_{l}(t)\|_{N-2}$
@@
where the last inequality is due to (2.16), (3.50), (3.51), the bounds $l<l_{\rm temp}$ and $l_{\rm temp}\|u_{q}\|_{1}\leq 1$, as well as the inequality $\|R_{l}(t)\|_{N}\lesssim l^{-N}\lambda_{q}^{-\gamma}\delta_{q+1}$ valid for $N\geq 0$.
@@
(4.26) $\|R_{T}^{\prime}(t)\|_{N}\lesssim\lambda_{q+1}^{N}\rho^{1/2}\left(\frac{\rho^{1/2}\mu^{-1}}{\lambda_{q+1}}+l\lambda_{q}\delta_{q}^{1/2}\right).$
@@
Therefore, the desired bounds follow from Lemmas˜3.8, 3.3 and 3.5, as well as
@@
4.1.2. Proof of Proposition˜4.1: Conclusion
We are ready to show Proposition˜4.1. First, bounds (4.11) and (4.12) follow directly from previous Lemmas˜4.2, 4.3 and 4.4, after some some standard computations. Instead, the bounds (4.9) and (4.10) are due to Lemmas˜4.2, 4.3 and 4.4 and Proposition˜3.3 combined with
(4.39) $\|Xw(t)\|_{N}\lesssim\sum_{j=0}^{N}\|X(t)\|_{j}\|w(t)\|_{N-j}$
applied with either $X(t):=R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}$ or $X(t):=D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right).$
@@ 5. Construction of the new current
(5.8) $+\left[(w_{o}\otimes w_{o}-c_{0})-\mathcal{R}(\mathord{{\rm div}}(w_{o}\otimes w_{o}-c_{0})+D_{t,l}w+w\cdot\nabla u_{l})\right]:\nabla(v_{l}+z_{q})^{T}-R_{q+1}:\nabla\overline{z}^{T}$
@@
(5.10) $-(z_{q+1}-z_{l})\cdot\mathord{{\rm div}}(w\otimes v_{l})+(v_{q}-v_{l})\cdot\mathord{{\rm div}}(w\otimes\overline{z})+w\cdot\mathord{{\rm div}}\Big\{(p_{q+1}-p_{l})\mathrm{Id}+R_{l}$
@@
(5.20) $f_{q+1}(t)$ $:=f_{A}(t)+f_{D_{2}}(t)+f_{S_{2}}(t):=\int_{0}^{t}\fint_{\mathbb{T}^{3}}\left(\varphi_{A}(x,s)+\varphi_{D_{2}}(x,s)+\varphi_{S_{2}}(x,s)\right)\,dx\,ds,$
@@ 5.1. Estimates on the current
Our main proposition on the current is the following Proposition˜5.1. The same caveat as before Proposition˜3.3 applies also here: the proof of this results could be skipped at a first reading. Roughly speaking, taking either a space derivative or an advective derivative of the new current $\phi_{q+1}$ “costs” a factor $\lambda_{q+1}$. On the other hand, the scalar function $f$ and its time derivative go to zero as $\lambda_{q+1}^{-\gamma}\delta_{q+2}$ and $\lambda_{q+1}^{-\gamma}\delta_{q+2}\delta_{q+1}^{1/2}$, respectively.
@@
5.1.1. Proof of Proposition˜5.1: bounds on $\phi_{A}$
@@
The result is due to Propositions˜A.3 and A.4 applied to $F=\phi_{A}$. Indeed, recall the decomposition of $\phi_{A}$ given by (5.21). Following the notation of Proposition˜A.3, let us define $a_{m,k}:=a^{1}_{m,k}+a^{2}_{m,k}+a^{3}_{m,k}$ where
@@
We intend to apply Proposition˜A.3 with $\lambda=\lambda_{q+1}$, a sequence
@@
Let us move to the estimate on $\|D_{t,q+1}\phi_{A}(t)\|_{N-1}$. We apply Proposition˜A.4 with $\lambda,s_{j}$ as above and $c=\epsilon$. In this case, we check the assumptions on $D_{t,l}a_{m,k}^{1}$ by the inductive estimates, Lemma˜3.8, the bound (3.64), Lemma˜3.4, and ˜3.5. In a similar fashion, in order to control the assumptions on $D_{t,l}a_{m,k}^{2}$ we use Bernstein inequality, (3.61) and [DLK23, Lemma A.4] to control, on one hand:
@@
Finally, as for the terms $D_{t,l}a_{m,k}^{3}$, we have by Lemma˜3.8
@@
At last, let us consider $|f_{A}(t)|+|f^{\prime}_{A}(t)|$. Due to the spatial and temporal partitions of unity, arguing as in Proposition˜3.3 it is sufficient to estimate $f_{m,k}(t):=\int_{0}^{t}\int_{\mathbb{T}^{3}}a_{m,k}(x,s)e^{i\lambda_{q+1}\xi_{m}(x,s)\cdot k}\,dx\,ds$. Then, [DLK23, Lemma A.2] implies that
(5.59) $|f_{A}(t)|+|f^{\prime}_{A}(t)|\lesssim\lambda_{q+1}^{-n_{0}+1}(s_{n_{0}+1}+s_{0}\,l^{-n_{0}-1})\lesssim\frac{\delta_{q+1}}{\epsilon}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
5.1.2. Proof of Proposition˜5.1: bounds on $\phi_{D}$
@@
Moving to the other terms, we have by (2.14), Lemma˜3.8, (3.90) and the bound $\mu^{-1}<\lambda_{q+1}$ that
@@
where in the last line we have used (5.66) above, Lemma˜3.8, the bounds (3.61), (3.51), and Proposition˜3.3. By similar computations, we can control also
@@
The next term in the decomposition (5.30) we deal with is $\phi_{q}-\phi_{l}$. In this case, we can repeat the computations reported in Lemma˜4.2, obtaining
(5.73) $\|\phi_{q}-\phi_{l}\|_{N}$ $\lesssim l^{1-N}\lambda_{q}^{1-\gamma}\delta_{q+1}^{3/2}(l^{-1}\delta^{1/2}_{q}l_{\rm temp}+1),$
@@
The cubic term in the velocity perturbation is controlled by means of Proposition˜3.3 as follows:
@@
(5.78) $=-\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}c_{m,k}e^{i\lambda_{q+1}k\cdot\xi_{m}}-R_{A}-R_{S}-w\odot(v_{q}-v_{l}),$
and then we can control it by Lemmas˜3.8, 4.3 and 4.4 as follows
@@
whereas the advective derivative can be controlled by using the aforementioned lemmas, Proposition˜3.3, and the estimates (3.51), (3.61)
@@
(5.81) $\quad+\left\|(v_{q}-v_{l})\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\left(D_{t,l}(c_{m,k})e^{i\lambda_{q+1}k\cdot\xi_{m}}+(u_{q+1}-u_{l})\cdot\nabla(c_{m,k}e^{i\lambda_{q+1}k\cdot\xi_{m}})\right)\right\|_{N-1}$
@@
Now, recalling the definition on the antidivergence stress $R_{A}$ given by (4.5), we observe that each $G_{i}$ has already been treated in the proof of Lemma˜4.3. Therefore, the estimates therein and (3.90) imply that
@@
Now, let us turn to Eq.˜5.85. Denote $H_{i}:=\mathcal{R}\mathbf{P}_{\geq 3\lambda/8}G_{i}:\nabla u_{l}^{T},$ then
(5.95) $\|D_{t,l}\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$ $\leq\|\mathcal{R}\left(\mathbf{P}_{\geq\lambda_{q+1}}D_{t,l}H_{i}+[u_{l}\cdot\nabla,\mathbf{P}_{\geq\lambda_{q+1}}]H_{i}\right)\|_{N-1}$
(5.96) $\quad+\|[u_{l}\cdot\nabla,\mathcal{R}]\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$
(5.97) $\lesssim\frac{1}{\lambda_{q+1}}\|D_{t,l}H_{i}+[u_{l}\cdot\nabla,\mathbf{P}_{\geq\lambda_{q+1}}]H_{i}\|_{N-1}+\|[u_{l}\cdot\nabla,\mathcal{R}]\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$

### v3

@@ 3.5. Estimates on the velocity perturbation
3.5.1. Proof of Proposition 3.1: Preliminary estimates
We report some estimates that will be helpful in the proof of Proposition 3.1, as well as in the remainder of the paper.
@@
Lemma 3.3 and (2.29) imply the following estimates, that will be useful later on.
@@
To keep the notation light we drop the time $t$ below, keeping in mind that all estimates are uniform in $t\leq\mathfrak{t}$. Moving to the proof of the lemma, notice that we have
@@
The same computations and Lemma 3.3 yield the result for $\|D_{t,l}(v_{q}-v_{l})\|_{N}$. ∎
@@
where $R_{comm}:=u_{l}\otimes u_{l}-\mathbf{P}_{\leq l^{-1}}(u_{q}\otimes u_{q})$ and $(\nabla u_{l})^{2}$ denotes the standard product between matrices $(\nabla u_{l})^{2}]^{i,j}=\sum_{k}\partial_{i}u^{k}\partial_{k}u^{j}$, then by [DLK23, Lemma A.3], Lemma 3.3, and the bound $L\,i_{q}^{-1/2-\delta}\leq\lambda^{2}_{q}\delta_{q}$ we have:
(3.65) $\|D_{t,l}\nabla u_{l}\|_{N}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}+L\,i_{q}^{-1/2-\delta}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}.$
Then, the estimate (3.63) on $\|D_{t,l}\nabla v_{l}\|_{N}$ follows from the identity $D_{t,l}\nabla v_{l}=D_{t,l}\nabla u_{l}-D_{t,l}\nabla z_{l}$ and computations similar to those above. ∎
3.5.2. Proof of Proposition 3.1: Estimates on the oscillatory and compressible Fourier coefficients
@@
Let us start with estimates aimed to control the Fourier coefficients appearing in $w$ and in the products $w_{o}\otimes w_{o}$ and $|w_{o}|^{2}w_{o}$, cf. Equations (3.23), (3.24), (3.25), (3.32).
@@
For $s\in\{0,1,2\}$ and $N\leq n_{0}+4$ , it holds for all $t\leq\mathfrak{t}$ that
@@
(3.70) $X_{I}(t,x):=\theta_{I}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(t,x)\overline{f}_{I}(t,x)\mathring{b}_{I,k}.$
@@
(3.71) $|D_{t,l}^{s}\theta_{I}(t)|\lesssim\epsilon^{-s}.$
Second, by Proposition A.1 and (3.63), together with the assumed bound $\epsilon\|u_{l}(t)\|_{1}<1$ and the algebraic identities
@@
(3.73) $\|D_{t,l}^{s}\overline{f_{I}}\|_{K}\lesssim l^{-K}(\lambda_{q}\delta_{q}^{1/2})^{s},\quad\text{for }s\in\{0,1,2\}.$
Third, by Proposition A.1, (3.18) and [BD15+, Proposition C.1], we have
@@
For $s=1$ instead, we have by (3.72), Proposition A.1 and (3.35):
@@
where $M_{m,n}$ is given by (3.43). Then, (3.72), (3.35), (3.83) imply that
(3.85) $\left\|D_{t,l}^{s}\left((\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\right\|_{N}\lesssim\epsilon^{-s}\mu^{-N}.$
Let us stress that we have exploited the bound (3.83) since $M_{m,n}$ is given by a finite summation of $X_{J}$ for $J\in\mathcal{I}_{\phi}$. Therefore,
@@
3.5.3. Proof of Proposition 3.1: Conclusion
@@
(3.88) $D_{t,l}^{s}w_{o}(t,x)=\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}D_{t,l}^{s}b_{m,k}(t,x)e^{i\lambda_{q+1}\xi_{m}(t,x)\cdot k},$
@@
(3.89) $\|e^{i\lambda_{q+1}\xi_{m}\cdot k}\|_{N}\lesssim|k|^{N}\lambda^{N}_{q+1},$
and thus the desired inequality follows from Lemma 3.6 and the analog of [DLK23, Equation (3.24)].
@@ 4. Construction of the new Reynolds stress and pressure
(4.8) $\operatorname{Tr}\big(R_{q+1}\big)=\operatorname{Tr}\big(R_{q}-R_{l}\big)+|w_{c}|^{2}+2w_{c}\cdot w_{o}+2(v_{q}-v_{l})\cdot w+2f.$
@@ 4.1. Estimates on the Reynolds stress
We denote $D_{t,q+1}:=\partial_{t}+u_{q+1}\cdot\nabla$. Our main proposition on the Reynolds stress is the following: morally speaking, taking either a space derivative or an advective derivative of the new Reynolds stress $R_{q+1}$ “costs” a factor $\lambda_{q+1}$.
@@
(4.10) $\left\|D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right)w(t)\right\|_{N-2}$ $\leq\lambda_{q+1}^{N-1-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2},$
@@
(4.12) $\left\|D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\,\right)\right\|_{N-2}$ $\leq\lambda_{q+1}^{N-1-2\gamma}\delta_{q+2}\delta_{q+1}^{1/2}.$
Remark 4.1.
The parameters $n_{0},h_{0}$ of previous proposition are given by:
@@
They will play a crucial role when perfoming estimates in which we need to apply the antidivergence operator $\mathcal{R}$, more specifically in the bounds (A.19), (A.46), (A.61) from the Appendix. The values of $n_{0},h_{0}$ ensure that
@@
4.1.1. Proof of Proposition 4.1: bounds on $R_{T},R_{A},R_{S}$.
As a first step towards proving Proposition 4.1, we separately give bounds on the stress tensors $R_{T},R_{A},R_{S}$ appearing in the decomposition of the new Reynolds stress $R_{q+1}$.
@@
(4.20) $(\mathbf{P}_{\leq l^{-1}}R_{q}-R_{l})(t,x)=\int_{-l_{\rm temp}}^{0}ds\,\eta_{l_{\rm temp}}(s)\int_{s}^{0}(D_{t,l}\mathbf{P}_{\leq l^{-1}}R_{q})(t+r,\mu(t;t+r,x))dr.$
@@
(4.23) $\|\mathbf{P}_{\leq l^{-1}}R_{q}-R_{l}\|_{N}\lesssim l^{-N}\lambda_{q}^{1-\gamma}\delta_{q+1}\delta_{q}^{1/2}l_{\rm temp}.$
@@
(4.25) $\|D_{t,q+1}(R_{q}-R_{l})\|_{N-2}$ $\leq\|D_{t,q+1}R_{q}\|_{N-2}+\|D_{t,q+1}R_{l}\|_{N-2}$
$\leq\|D_{t,q}R_{q}\|_{N-2}+\|(u_{q+1}-u_{q})\cdot\nabla R_{q}\|_{N-2}$
$+\|D_{t,l}R_{l}\|_{N-2}+\|(u_{q+1}-u_{l})\cdot\nabla R_{l}\|_{N-2}$
@@
where the last inequality is due to (2.16), (3.49), (3.50), the bounds $l<l_{\rm temp}$ and $l_{\rm temp}\|u_{q}\|_{1}\leq 1$, as well as the inequality $\|R_{l}\|_{N}\lesssim l^{-N}\lambda_{q}^{-\gamma}\delta_{q+1}$ valid for $N\geq 0$.
@@
(4.26) $\|R_{T}^{\prime}\|_{N}\lesssim\lambda_{q+1}^{N}\rho^{1/2}\left(\frac{\rho^{1/2}\mu^{-1}}{\lambda_{q+1}}+l\lambda_{q}\delta_{q}^{1/2}\right).$
@@
Therefore, the desired bounds follow from Lemma 3.6,Proposition 3.1, Corollary 3.1, as well as
@@
4.1.2. Proof of Proposition 4.1: Conclusion
We are ready to show Proposition 4.1. First, bounds (4.11) and (4.12) follow directly from previous Lemma 4.1, Lemma 4.2, Lemma 4.3, after some some standard computations. Instead, the bounds (4.9) and (4.10) are due to Lemma 4.1, Lemma 4.2, Lemma 4.3 and Proposition 3.1 combined with
(4.39) $\|Xw\|_{N}\lesssim\sum_{j=0}^{N}\|X\|_{j}\|w\|_{N-j}$
applied with either $X(t,x):=R_{q+1}(t,x)-\frac{2}{3}f(t)\mathrm{Id}$ or $X(t,x):=D_{t,q+1}\left(R_{q+1}(t,x)-\frac{2}{3}f(t)\mathrm{Id}\right).$
@@ 5. Construction of the new current
(5.8) $+\left((w_{o}\otimes w_{o}-c_{0})-R_{A}\right):\nabla(v_{l}+z_{q})^{T}-R_{q+1}:\nabla\overline{z}^{T}$
@@
(5.10) $-(z_{q+1}-z_{l})\cdot\mathord{{\rm div}}(w\otimes v_{l})+(v_{q}-v_{l})\cdot\mathord{{\rm div}}(w\otimes\overline{z})+w\cdot\mathord{{\rm div}}\Big\{(p_{q+1}-p_{l})\mathrm{Id}+\mathbf{P}_{\leq l^{-1}}R_{q}$
@@
(5.20) $f_{q+1}(t)$ $:=f_{A}(t)+f_{D_{2}}(t)+f_{S_{2}}(t):=\int_{0}^{t}\fint_{\mathbb{T}^{3}}\left(\varphi_{A}(s,x)+\varphi_{D_{2}}(s,x)+\varphi_{S_{2}}(s,x)\right)\,dx\,ds,$
@@ 5.1. Estimates on the current
Our main proposition on the current is the following Proposition 5.1. The same caveat as before Proposition 3.1 applies also here: the proof of this results could be skipped at a first reading. Roughly speaking, taking either a space derivative or an advective derivative of the new current $\phi_{q+1}$ “costs” a factor $\lambda_{q+1}$. On the other hand, the scalar function $f$ and its time derivative go to zero as $\lambda_{q+1}^{-\gamma}\delta_{q+2}$ and $\lambda_{q+1}^{-\gamma}\delta_{q+2}\delta_{q+1}^{1/2}$, respectively.
@@
5.1.1. Proof of Proposition 5.1: bounds on $\phi_{A}$
@@
The result is due to Proposition A.2, Proposition A.3 applied to $F=\varphi_{A}$. Indeed, recall the decomposition of $\varphi_{A}$ given by (5.22). Following the notation of Proposition A.2, let us define $a_{m,k}:=a^{1}_{m,k}+a^{2}_{m,k}+a^{3}_{m,k}$ where
@@
We intend to apply Proposition A.2 with $\lambda=\lambda_{q+1}$, a sequence
@@
Let us move to the estimate on $\|D_{t,q+1}\phi_{A}\|_{N-1}$. We apply Proposition A.3 with $\lambda,s_{j}$ as above and $c=\epsilon$. In this case, we check the assumptions on $D_{t,l}a_{m,k}^{1}$ by the inductive estimates, Lemma 3.6, the bound (3.63), Lemma 3.3, and Corollary 3.1. In a similar fashion, in order to control the assumptions on $D_{t,l}a_{m,k}^{2}$ we use Bernstein inequality, (3.60) and [DLK23, Lemma A.4] to control, on one hand:
@@
Finally, as for the terms $D_{t,l}a_{m,k}^{3}$, we have by Lemma 3.6
@@
At last, let us consider $|f_{A}|+|f^{\prime}_{A}|$. Due to the spatial and temporal partitions of unity, arguing as in Proposition 3.1 it is sufficient to estimate $f_{m,k}(t):=\int_{0}^{t}\int_{\mathbb{T}^{3}}a_{m,k}(s,x)e^{i\lambda_{q+1}\xi_{m}(s,x)\cdot k}\,dx\,ds$. Then, [DLK23, Lemma A.2] implies that
(5.59) $|f_{A}|+|f^{\prime}_{A}|\lesssim\lambda_{q+1}^{-n_{0}+1}(s_{n_{0}+1}+s_{0}\,l^{-n_{0}-1})\lesssim\frac{\delta_{q+1}}{\epsilon}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
5.1.2. Proof of Proposition 5.1: bounds on $\phi_{D}$
@@
Moving to the other terms, we have by (2.14), Lemma 3.6, (3.89) and the bound $\mu^{-1}<\lambda_{q+1}$ that
@@
where in the last line we have used (5.66) above, Lemma 3.6, the bounds (3.60), (3.50), and Proposition 3.1. By similar computations, we can control also
@@
The next term in the decomposition (5.30) we deal with is $\phi_{q}-\phi_{l}$. In this case, we can repeat the computations reported in Lemma 4.1, obtaining
(5.73) $\|\phi_{q}-\phi_{l}\|_{N}$ $\lesssim l^{1-N}\lambda_{q}^{1-\gamma}\delta_{q+1}^{3/2}(l^{-1}\delta_{q}^{1/2}l_{\rm temp}+1),$
@@
The cubic term in the velocity perturbation is controlled by means of Proposition 3.1 as follows:
@@
(5.78) $=-\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}c_{m,k}e^{i\lambda_{q+1}\xi_{m}\cdot k}-R_{A}-R_{S}-w\odot(v_{q}-v_{l}),$
and then we can control it by Lemma 3.6, Lemma 4.2, Lemma 4.3 as follows
@@
whereas the advective derivative can be controlled by using the aforementioned lemmas, Proposition 3.1, and the estimates (3.50), (3.60)
@@
(5.81) $\quad+\left\|(v_{q}-v_{l})\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\left(D_{t,l}(c_{m,k})e^{i\lambda_{q+1}\xi_{m}\cdot k}+(u_{q+1}-u_{l})\cdot\nabla(c_{m,k}e^{i\lambda_{q+1}\xi_{m}\cdot k})\right)\right\|_{N-1}$
@@
Now, recalling the definition on the antidivergence stress $R_{A}$ given by (4.5), we observe that each $G_{i}$ has already been treated in the proof of Lemma 4.2. Therefore, the estimates therein and (3.89) imply that
@@
Now, let us turn to (5.85). Denote $H_{i}:=\mathcal{R}\mathbf{P}_{\geq 3\lambda/8}G_{i}:\nabla u_{l}^{T},$ then
(5.95) $\|D_{t,l}\overline{\mathcal{R}}\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$ $\leq\|\overline{\mathcal{R}}\left(\mathbf{P}_{\geq\lambda_{q+1}}D_{t,l}H_{i}+[u_{l}\cdot\nabla,\mathbf{P}_{\geq\lambda_{q+1}}]H_{i}\right)\|_{N-1}$
(5.96) $\quad+\|[u_{l}\cdot\nabla,\overline{\mathcal{R}}]\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$
(5.97) $\lesssim\frac{1}{\lambda_{q+1}}\|D_{t,l}H_{i}+[u_{l}\cdot\nabla,\mathbf{P}_{\geq\lambda_{q+1}}]H_{i}\|_{N-1}+\|[u_{l}\cdot\nabla,\overline{\mathcal{R}}]\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$